OpenStudy (anonymous):

Mrs. Walters drove 360 miles on 25 gallons of gas. Mr. Walters drove 480 miles on 32 gallons of gas. Find the mileage per gallon for each car. Which car is getting the better mileage per gallon?

OpenStudy (texaschic101):

Mrs. Walters... 360/25 = x/1 (360 miles to 25 gallons = x miles to 1 gallon)- cross multiply (25)(x) = (360)(1) 25x = 360 x = 360/25 x = 14 2/5 miles per gallon Mr. Walters ... 480/32 = x/1 (480 miles to 32 gallons = x miles to 1 gallon) - cross multiply (32)(x) = (480)(1) 32x = 480 x = 480/32 x = 15 miles per gallon Mr. Walters is getting better mileage per gallon.